OpenStudy (anonymous):

A triangle is graphed in the coordinate plane. The vertices of the triangle have coordinates (–3, 3), (5, 3), and (5, –3). What is the perimeter of the triangle? A. 22 B. 24 C. 28 D. 30

OpenStudy (amistre64):

your simply going to have to apply the distance formula 3 times around ....

OpenStudy (amistre64):

or, if they line up nicely, youll only have to use it once

OpenStudy (amistre64):

 yeah, plotting them shows that 2 sides are simple enough to assess. the third side is a pythagorean process.

OpenStudy (anonymous):

Ok so what is the equation or solution that I use?

OpenStudy (amistre64):

well the 2 sides are simply counting .... the third side you can use the pythagorean thrm for: a^2 + b^2 = c^2, and solve for x

OpenStudy (amistre64):

*solve for c .... they had to put that key right next to the x :)

OpenStudy (anonymous):

Thank you

OpenStudy (amistre64):

good luck

OpenStudy (amistre64):

if you want to check your solution ... feel free to post it

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

x=c over b ??

OpenStudy (amistre64):

the 2 sides (a and b) are simply counting .... the third side you can use the pythagorean thrm for: a^2 + b^2 = c^2, and solve for c (the hypotenuse of a rt triangle) \[c=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

a2 + b2 = ??

OpenStudy (anonymous):

Or give the solution to the problem so I can solve for x.

OpenStudy (amistre64):

there is no x to solve ....

OpenStudy (amistre64):

what do you get for the 2 sides, a and b?